If we have 10 pulses occurring in 2 micro seconds what would be the PRF
1
Expert's answer
2019-07-23T15:55:31-0400

If we have 10 pulses occurring in 2 micro seconds, then in 1 second there would be "10\cdot \frac{1}{2 \cdot 10^{-6}}=5\cdot 10^6" pulses. So, the PRF would be 5 MHz.